Determine in each case whether the given planes are parallel or perpendicular

a) x + y + 3z 10 = 0 and x + 2y - z = 1

b) 3x - 2y + z - 6 = 0 and 4x + 2y - 4z = 0

c) 3x + y + z - 1 = 0 and -x + 2y + z + 3 = 0


1
Expert's answer
2021-07-01T17:15:39-0400

The direction ratio of plane Pi: aix+biy+ciy+ di=0 is <ai,bi,ci>.

Two planes P1 and P2 are parallel if

a1/a2 = b1/b2 = c1/c2 and perpendicular if a1a2 +b1 b2 + c1c2 =0


a). Here <a1,b1,c1>=<1,1,3>and <a2,b2,c3>=<1,2,-1>

a1/a2 = 1, b1/b2 =1/2, c1/c2=-3

a1a2 +b1 b2 + c1c2 =1+2-3=0

Hence these two planes are perpendicular.


b).Here <a1,b1,c1>=<3,-2,1>and <a2,b2,c3>=<4,2,-4>

a1/a2 = 3/4, b1/b2 =-1, c1/c2=-1/4

a1a2 +b1 b2 + c1c2 =12-4-4=4

Hence these two planes are neither parallell nor perpendicular.


c).Here <a1,b1,c1>=<3,1,1>and <a2,b2,c3>=<-1,2,1>

a1/a2 = -3, b1/b2 =1/2, c1/c2=1

a1a2 +b1 b2 + c1c2 =-3+2+1=0

Hence these two planes are perpendicular.